### v4

@@ 1 Introduction
where the constraint $x^{s}-\overline{x}=0,\ \forall s\in S$, enforces non-anticipativity for the first-stage decisions. The RDEM problem (4) could be fully decomposed into $\lvert S\rvert$ non-convex MIQCQP problems if one could remove the set of linear constraints $x^{s}-\overline{x}=0,\ \forall s\in S$, that relates variables from distinct sub-problems, a structure commonly known as complicating constraints.
@@ 2.1 Reformulated normalized multiparametric disaggregation technique (RNMDT)
The normalised multiparametric disaggregation technique (NMDT) is an efficient technique to relax quadratic terms in non-convex MIQCQP problems (Castro, 2016a). NMDT involves the discretisation of the domain of one variable in each bi-linear term. The discretisation procedure in NMDT closely resembles the piecewise McCormick envelopes approach (McCormick, 1976) as it involves splitting the variable domain into a number of uniform partitions. The accuracy of the NMDT approximation of the primary non-convex MIQCIP problem is directly related to the size of these partitions and can be made arbitrarily precise. However, improving the accuracy of NMDT approximation follows the increase in the number of continuous and binary variables. Therefore, there is a trade-off between the accuracy of the NMDT approximation and its computational tractability.
@@ 3 Dual decomposition
where $\epsilon_{BB}>0$. The branching index $i$ is chosen based on the measure of the dispersion in the first-stage scenario solutions, e.g., if the dispersion of the component $i:\ \delta_{i}=\max_{s\in S}x_{N,i}^{*,s}-\min_{s\in S}x_{N,i}^{*,s}$ is zero, this should imply the non-anticipativity of this component
@@
Algorithm 4 summarises adaptation of the branch-and-bound method presented by Carøe and Schultz (1999) that hereafter we refer to as p-BnB. For each branch-and-bound node $N\in T$, we generate node sub-problem (9) and compute its dual bound value $z^{*}_{N}$ as well as corresponding optimal dual and primal variables values $(\mu_{N}^{*,s})_{s\in S}$ and $(x_{N}^{*,s},y_{N}^{*,s},w_{N}^{*,s})_{s\in S}$, respectively, by applying Algorithm 1 or 2. If the dual bound value $z^{*}_{N}>z_{UB}$, the node $N$ is fathomed. Otherwise, we check whether solution $x_{N}^{*,s}$ violates non-anticipativity or integrality conditions. If that is the case, we perform branching as described in (25)–(26) on the most fractional variable $\overline{x}_{N,i}$ if $x_{N}^{*,s}$ violates integrality conditions. Otherwise, we perform branching as described in (27)–(28) on the variable with the largest dispersion $\delta_{i}$ if $x_{N}^{*,s}$ violates non-anticipativity conditions. If $x_{N}^{*,s}$ satisfies both non-anticipativity and integrality conditions, we update the best upper bound value $z_{UB}=z^{*}_{N}$ and best solution value $x^{*}=\overline{x}_{N}$. Lastly, we update the best lower bound value $z_{LB}$ by setting it to the smallest dual bound value $z^{*}_{N}$ among the nodes $N$ that are yet to be fathomed. The algorithm continues until the set $T$ is empty.
@@ 4.1 Design of experiments
Table 2 presents the parameter values used in the experiments for the proximal BM (Algorithm (1)) and the FWPH (Algorithm 2). In addition to the parameters stated in Table 2, the maximum number of iterations for the proximal BM ($k_{max}$) was set to 1000. The maximum numbers of iterations for the FWPH algorithm ($k_{max}$) and simplicial decomposition method ($t_{max}$) were set to 1000 and 1, respectively. We considered $t_{max}=1$ as Boland et al. (2019) suggest that considering higher values of $t_{max}$ parameter commonly increases the computational burden while not guaranteeing an adequate increase in the quality of the approximation of $\text{conv}(G^{s})$ compared to the case when $t_{max}=1$. It is also worth mentioning that since $t_{max}=1$ for the simplicial decomposition method, the $\epsilon_{SDM}$ can be set to any arbitrary value as the condition $\Gamma^{t}\leq\epsilon_{SDM}$ is not going to be checked for the algorithm termination. Following (Boland et al., 2018, Proposition 3.3), since we assume $t_{max}=1$, to guarantee the convergence of Algorithm 2 one should generate initial sets $\{(V_{0}^{s})_{s\ inS}\}$ such that $\bigcap_{s\in S}\text{Proj}_{x}(\text{conv}(V_{0}^{s}))\neq\emptyset$. Therefore, to initialise $(V^{s}_{0})_{s\in S}$, we took one arbitrary scenario (in our case the first scenario in $S$, i.e., $s=1$) and set $V^{1}_{0}=\{(x^{1}_{0},y^{1}_{0},w^{1}_{0})\}$. Further, for each $s\in S,s\neq 1$, we initialised $V^{s}_{0}=\{{(x^{s}_{0},y^{s}_{0},w^{s}_{0}),(x^{1}_{0},\overline{y}^{s},\overline{w}^{s})}\}$, where $(x^{s}_{0},y^{s}_{0},w^{s}_{0})$ solves $L^{s}(\mu^{s}_{0})$ and $(\overline{y}^{s},\overline{w}^{s}))$ solves,
@@
Starting dual multipliers values $\mu_{0}$ for Algorithms 1 and 2 were set $\mu_{0}=0$. To set the first-stage variables $(x^{s}_{0})_{s\in S}$ for Algorithm 2, we considered the solution of the p-Lagrangian dual function (6) for a fixed value of the dual variable $\mu=\mu_{0}$. The tolerances $\epsilon_{BB}$ and $\epsilon_{NAC}$ for the p-BnB (Algorithm 4) were set to $10^{-6}$. As a time limit for solving each distinct instance, we considered one hour.
@@ 4.2 Numerical results
Interestingly, for the setting where the matrices are denser, FWPH outperforms BM as the Lagrangian dual solver in Instances 4 and 5, which is a consequence of FWPH leading the algorithm to explore less nodes than BM. In either case, the number of nodes explored is small. As can be seen in Table 5, the maximum number of nodes explored by p-BnB was only 11, for Instance 5 using the proximal BM, while the average number of nodes explored was five. This is because despite the very high density of the quadratic matrices (90%) in the instances, at the very first node, p-BnB was able to generate a solution with a tight dual bound, on average being 0.01%. In comparison, the average dual bound generated within one hour by solving the Instances 1-5 with Gurobi was 4.98%.

### v5

@@ 1 Introduction
where the constraint $x^{s}-\overline{x}=0,\ \forall s\in S$, enforces non-anticipativity for the first-stage decisions and $\overline{x}$ is an auxiliary variable used to enforce nonanticipativity for all $x^{s},\ \forall s\in S$. The RDEM problem (4) could be fully decomposed into $\lvert S\rvert$ non-convex MIQCQP problems if one could remove the set of linear constraints $x^{s}-\overline{x}=0,\ \forall s\in S$, that relates variables from distinct sub-problems, a structure commonly known as complicating constraints.
@@ 2.1 Reformulated normalized multiparametric disaggregation technique (RNMDT)
The normalised multiparametric disaggregation technique (NMDT) is an efficient technique to relax quadratic terms in non-convex MIQCQP problems (Castro, 2016a). NMDT involves the discretisation of the domain of one variable in each bi-linear term. The discretisation procedure in NMDT closely resembles the piecewise McCormick envelopes approach (Bergamini et al., 2005; McCormick, 1976) as it involves splitting the variable domain into a number of uniform partitions. The accuracy of the NMDT approximation of the primary non-convex MIQCIP problem is directly related to the size of these partitions and can be made arbitrarily precise. However, improving the accuracy of NMDT approximation follows the increase in the number of continuous and binary variables. Therefore, there is a trade-off between the accuracy of the NMDT approximation and its computational tractability.
@@ 3 Dual decomposition
There are multiple approaches for selecting the integer index $i$ of fractional-valued variable, i.e., $\lfloor\overline{x}_{N,i}\rfloor<\overline{x}_{N,i}<\lceil\overline{x}_{N,i}\rceil$ to perform branching. Morrison et al. (2016) describes several common methods, of which we highlight the following:
  • Branching on the most (least) fractional or most (least) infeasible variable (see, for example, Achterberg et al. (2005); Ortega and Wolsey (2003)): This approach involves selecting the integer index $i$ such that the fractional part of $\overline{x}_{N,i}$ is closest to (or furthest from) 0.5.
  • Pseudocost branching (see, for example, Bénichou et al. (1971)): In this method, the integer index $i$ is chosen based on the expected impact of branching on $\overline{x}_{N,i}$ on the objective function, leveraging the history of previous branching decisions in the tree.
  • Strong branching (see, for example, Achterberg et al. (2005); Achterberg (2007)): This technique recommends selecting the integer index $i$ such that branching on $\overline{x}_{N,i}$ results in the greatest change to the objective function. This is typically evaluated by solving LP relaxations for the child nodes generated by branching on $\overline{x}_{N,i}$.
In addition to these strategies, the authors in Morrison et al. (2016) highlight several other methods, including combining strong and pseudocost branching (see, for example, Linderoth and Savelsbergh (1999)), backdoor branching, which uses an auxiliary integer program to identify a subset of indices that reduce the search tree size (see, for example, Fischetti and Monaci (2011)), and information-theoretic branching, where indices are chosen to minimize uncertainty in the sub-problems (see, for example, Gilpin and Sandholm (2011)).
In our approach, a predominant feature is that we typically have a small search tree and that each node typically requires considerable computational work. The combination of these two characteristics led us to lean towards a simpler branching strategy.
@@
where $\epsilon_{BB}>0$. The branching index $i$ is chosen based on the measure of the dispersion in the first-stage scenario solutions, e.g., if the dispersion of the component $i:\ \sigma_{i}=\max_{s\in S}x_{N,i}^{*,s}-\min_{s\in S}x_{N,i}^{*,s}$ is zero, this should imply the non-anticipativity of this component
@@
The branch-and-bound search requires that any node not pruned or fathomed must be explored. Therefore, reducing the size of the search tree is directly tied to the efficiency of the pruning strategy. Morrison et al. (2016) highlight the following major groups of strategies
  • Pruning by bound: This strategy involves pruning nodes whose sub-problem objective value is worse (higher in the case of minimisation) than the incumbent solution objective value (see, for example, Arora et al. (2002); Vilà and Pereira (2014)). The bounds of the sub-problems can be calculated either by solving a linear relaxation of the node sub-problems or by using duality techniques.
  • Pruning by the dominance relations: this strategy involves pruning a node if another node’s sub-problem dominates its sub-problem. A sub-problem is said to be dominated if another sub-problem guarantees a better solution or an equally good solution that is less computationally intensive (see, for example, Sewell et al. (2012); Fischetti and Salvagnin (2010)). Dominance rules can be defined based on whether memory is used to store previously explored sub-problems. The memory-based dominance involves comparing unexplored nodes to sub-problems of already generated and stored nodes. If a stored sub-problem dominates a new one, the new sub-problem can be pruned. The non-memory-based dominance involves determining the existence of a dominating node without relying on stored sub-problems. It checks whether a dominance relation holds regardless of whether the dominating node has been explored.
In the proposed approach, we prune nodes based on the lower bounds generated by solving the dual sub-problems, for the same reasons that we chose simpler branching rules (i.e., small search trees with computationally expensive nodes).
Algorithm 4 summarises adaptation of the branch-and-bound method presented by Carøe and Schultz (1999) that hereafter we refer to as p-BnB. For each branch-and-bound node $N\in T$, we generate node sub-problem (9) and compute its dual bound value $z^{*}_{N}$ as well as corresponding optimal dual and primal variables values $(\mu_{N}^{*,s})_{s\in S}$ and $(x_{N}^{*,s},y_{N}^{*,s},w_{N}^{*,s})_{s\in S}$, respectively, by applying Algorithm 1 or 2. If the dual bound value $z^{*}_{N}>z_{UB}$ or one of the $N$ sub-problems ${s\in S}$ is infeasible, the node $N$ is fathomed. Otherwise, we check whether solution $x_{N}^{*,s}$ violates non-anticipativity or integrality conditions. If that is the case, we perform branching as described in (25)–(26) on the most fractional variable $\overline{x}_{N,i}$ if $x_{N}^{*,s}$ violates integrality conditions. Otherwise, we perform branching as described in (27)–(28) on the variable with the largest dispersion $\sigma_{i}$ if $x_{N}^{*,s}$ violates non-anticipativity conditions. If $x_{N}^{*,s}$ satisfies both non-anticipativity and integrality conditions, we update the best upper bound value $z_{UB}=z^{*}_{N}$ and best solution value $x^{*}=\overline{x}_{N}$. Lastly, we update the best lower bound value $z_{LB}$ by setting it to the smallest dual bound value $z^{*}_{N}$ among the nodes $N$ that are yet to be fathomed. The algorithm continues until the set $T$ is empty.
@@
To clarify how the elements composing Algorithm 4 interact, we provide a graphical representation of the algorithm’s structure in Figure 1. Additionally, Section C presents an illustrative example, showcasing the procedures described in Algorithm 4 and demonstrating how they apply to a small problem instance.
Figure 1: Graphical representation of Algorithm 4 structure
@@ 4.1 Design of experiments
Table 2 presents the parameter values used in the experiments for the proximal BM (Algorithm 1) and the FWPH (Algorithm 2). In addition to the parameters stated in Table 2, the maximum number of iterations for the proximal BM ($k_{max}$) was set to 1000. The maximum numbers of iterations for the FWPH algorithm ($k_{max}$) and simplicial decomposition method ($t_{max}$) were set to 1000 and 1, respectively. We considered $t_{max}=1$ as Boland et al. (2019) suggest that considering higher values of $t_{max}$ parameter commonly increases the computational burden while not guaranteeing an adequate increase in the quality of the approximation of $\text{conv}(G^{s})$ compared to the case when $t_{max}=1$. It is also worth mentioning that since $t_{max}=1$ for the simplicial decomposition method, the $\epsilon_{SDM}$ can be set to any arbitrary value as the condition $\Gamma^{t}\leq\epsilon_{SDM}$ is not going to be checked for the algorithm termination. Following (Boland et al., 2018, Proposition 3.3), since we assume $t_{max}=1$, to guarantee the convergence of Algorithm 2 one should generate initial sets $\{(V_{0}^{s})_{s\ inS}\}$ such that $\bigcap_{s\in S}\text{Proj}_{x}(\text{conv}(V_{0}^{s}))\neq\emptyset$. Therefore, to initialise $(V^{s}_{0})_{s\in S}$, we took one arbitrary scenario (in our case the first scenario in $S$, i.e., $s=1$) and set $V^{1}_{0}=\{(x^{1}_{0},y^{1}_{0},w^{1}_{0})\}$. Further, for each $s\in S,s\neq 1$, we initialised $V^{s}_{0}=\{{(x^{s}_{0},y^{s}_{0},w^{s}_{0}),(x^{1}_{0},\overline{y}^{s},\overline{w}^{s})}\}$, where $(x^{s}_{0},y^{s}_{0},w^{s}_{0})$ solves $L^{s}(\mu^{s}_{0})$ and $(\overline{y}^{s},\overline{w}^{s}))$ solves,
@@
Starting dual multipliers values $\mu_{0}$ for Algorithms 1 and 2 were set $\mu_{0}=0$. To set the first-stage variables $(x^{s}_{0})_{s\in S}$ for Algorithm 2, we considered the solution of the p-Lagrangian dual function (6) for a fixed value of the dual variable $\mu=\mu_{0}$. The tolerances $\epsilon_{BB}$ and $\epsilon_{NAC}$ for the p-BnB (Algorithm 4) were set to $10^{-6}$. As a time limit for solving each distinct instance, we considered one hour. In case multiple integer indices $i_{1},\dots,i_{Int}$ are available for branching on the variable $\overline{x}_{N,i}$ in p-BnB, we simply choose the first index $i_{1}$.
@@ 4.2 Numerical results
Interestingly, for the setting where the matrices are denser, FWPH outperforms BM as the Lagrangian dual solver in Instances 4 and 5, which is a consequence of FWPH leading the algorithm to explore less nodes than BM. In either case, the number of nodes explored is small. As can be seen in Table 5, the maximum number of nodes explored by p-BnB was only 11, for Instance 5 using the proximal BM, while the average number of nodes explored was five. This is because despite the very high density of the quadratic matrices (90%) in the instances, at the very first node, p-BnB was able to generate a solution with a tight dual bound, on average being 0.01%. In comparison, the average dual bound generated within one hour by solving the Instances 1-5 with Gurobi was 4.98%. Further details on each individual Instance are provided in Appendix B.
@@ Appendix A Appendix A
Appendix B Appendix B
Table 7: Average dual bound generated within one hour by solving the Instances 1-5 with Gurobi solver
Instance dual bound (%)
1 0.00
2 4.71
3 6.52
4 8.54
5 5.11
Appendix C Appendix C
This section presents an illustrative example showcasing the application of Algorithm 4. Consider Problem (4) with 2 scenarios $S=\{s_{1},s_{2}\}$, two first-stage variables $x_{1}$ and $x_{2}$, one second-stage variable $y$, and one constraint defined per scenario. Then, Problem (4) can be formulated as follows:
$z^{\text{SMIP}}=\operatorname*{min.}_{x_{1}^{1},x_{1}^{2},x_{2}^{1},x_{2}^{2},y^{1},y^{2}}\leavevmode\nobreak$ $-16.8y^{1}y^{1}-9.8y^{1}$ (29)
$-4.5y^{2}y^{2}-7.9y^{2}$ (30)
$-53.2x_{1}^{1}-23.2x_{2}^{1}-9.0x_{1}^{2}-3.9x_{2}^{2}$ (31)
s.t.:
$78.1x_{1}^{1}+67.0x_{2}^{1}+16.8y^{1}\leq 57108.7$ (32)
$77.0y^{2}y^{2}+45.3x_{1}^{2}+30.2x_{2}^{2}+0.1y^{2}\leq 56702.4$ (33)
$x_{1}^{1}-\hat{x}_{1}=0$ (34)
$x_{1}^{2}-\hat{x}_{1}=0$ (35)
$x_{2}^{1}-\hat{x}_{2}=0$ (36)
$x_{2}^{2}-\hat{x}_{2}=0$ (37)
$0.0\leq x_{1}^{1},x_{1}^{2}\leq 5.0$ (38)
$0.0\leq x_{2}^{1},x_{2}^{2}\leq 7.0$ (39)
$0.0\leq y^{1}\leq 9.0$ (40)
$0.0\leq y^{2}\leq 8.0$ (41)
$x_{1}^{1},\ x_{1}^{2},\ x_{2}^{1},\ x_{2}^{2}\ \in\ \mathbb{Z}$ (42)
Where, (34)- (37) are non-anticipativity conditions.
Applying p-Lagrangian relaxation we get Problem (6) with two sub-problems (44) and (45)
$L^{1}(\mu^{1})=\operatorname*{min.}_{\omega^{1}}\leavevmode\nobreak$ $\frac{1}{\pi^{1}}\left(-16.8w^{1}-53.3x_{1}^{1}-23.2x_{2}^{1}-9.8y^{1}+\mu^{1}_{1}(x_{1}^{1}-\hat{x}_{1})+\mu^{1}_{2}(x_{2}^{1}-\hat{x}_{2})\right)$ (44)
s.t.:
$y^{1}-9\Delta y^{1}-4.5z_{-1}^{1}=0$
$-4.5\hat{y}_{-1}^{1}+w^{1}-9\Delta w^{1}=0$
$78.1x_{1}^{1}+67.0x_{2}^{1}+16.8y^{1}\leq 57108.7$
$0.5y^{1}+9\Delta y^{1}-\Delta w^{1}\leq 4.5$
$-0.5y^{1}+\Delta w^{1}\leq 0$
$-\Delta w^{1}\leq 0$
$-9\Delta y^{1}+\Delta w^{1}\leq 0$
$-\hat{y}_{-1}^{1}\leq 0$
$\hat{y}_{-1}^{1}-9z_{-1}^{1}\leq 0$
$-y^{1}+\hat{y}_{-1}^{1}\leq 0$
$y^{1}-\hat{y}_{-1}^{1}+9z_{-1}^{1}\leq 9.0$
$0.0\leq x_{1}^{1}\leq 5.0$
$0.0\leq x_{2}^{1}\leq 7.0$
$0.0\leq y^{1}\leq 9.0$
$0.0\leq\Delta y^{1}\leq 0.5$
$x_{1}^{1},x_{2}^{1}\in\mathbb{Z}$
$z_{-1}^{1}\in\{0,1\}$
$L^{2}(\mu^{2})=\operatorname*{min.}_{\omega^{2}}\leavevmode\nobreak$ $\frac{1}{\pi^{2}}\left(-4.5w^{2}-9.0x_{1}^{2}-3.9x_{2}^{2}-7.9y^{2}+\mu^{2}_{1}(x_{1}^{2}-\hat{x}_{1})+\mu^{2}_{1}(x_{2}^{2}-\hat{x}_{2})\right)$ (45)
s.t.:
$y^{2}-8\Delta y^{2}-4z_{-1}^{2}=0$
$-4\hat{y}_{-1}^{2}+w^{2}-8\Delta w^{2}=0$
$45.3x_{1}^{2}+30.2x_{2}^{2}+0.1y^{2}+77w^{2}\leq 56702.4$
$0.5y^{2}+8\Delta y^{2}-\Delta w^{2}\leq 4.0$
$-0.5y^{2}+\Delta w^{2}\leq 0$
$-\Delta w^{2}\leq 0$
$-8\Delta y^{2}+\Delta w^{2}\leq 0$
$-\hat{y}_{-1}^{2}\leq 0$
$\hat{y}_{-1}^{2}-8z_{-1}^{2}\leq 0$
$-y^{2}+\hat{y}_{-1}^{2}\leq 0$
$y^{2}-\hat{y}_{-1}^{2}+8z_{-1}^{2}\leq 8.0$
$0.0\leq x_{1}^{2}\leq 5.0$
$0.0\leq x_{2}^{2}\leq 7.0$
$0.0\leq y^{2}\leq 8.0$
$0.0\leq\Delta y^{2}\leq 0.5$
$x_{1}^{2},x_{2}^{2}\in\mathbb{Z}$
$z_{-1}^{2}\in\{0,1\}$
The terms $\pi^{1}$ and $\pi^{2}$ correspond to the probabilities of Problems (44) and (45), respectively, and
$\omega^{1}=\{x_{1}^{1},x_{2}^{1},\hat{x_{1}},\hat{x_{2}},y^{1},\Delta y^{1},\Delta w^{1},w^{1},\hat{y}_{-1}^{1},z_{-1}^{1}\},$
and
$\omega^{2}=\{x_{1}^{2},x_{2}^{2},\hat{x_{1}},\hat{x_{2}},y^{2},\Delta y^{2},\Delta w^{2},w^{2},\hat{y}_{-1}^{2},z_{-1}^{2}\}.$
Then the first root node of the Algorithm 4 will be
$z_{LD}=\operatorname*{max.}_{\mu^{1},\mu^{2}}\left\{\pi^{1}L^{1}(\mu^{1})+\pi^{2}L^{2}(\mu^{2}):\pi^{1}\mu^{1}+\pi^{2}\mu^{2}=0\right\}.$ (46)
Applying Algorithm 1 to solve Problem (46) yields the solution in just one iteration, due to the simple structure and small scale of the problem. A similar result occurs when applying Algorithm 2, which also converges to the solution $z_{LB}$ in one iteration. The solution is as follows:
Table 8: Result of applying Algorithm 1 and Algorithm 2 to Problem (46)
Number of iterations 1
Objective function value $z_{LB}=-2299.62$
First-stage variables’ values $x_{1}^{1}=5.0$, $x_{2}^{1}=7.0$, $x_{1}^{2}=5.0$, $x_{2}^{2}=7.0$
Second-stage variables’ values $y^{1}=9.0$, $y^{2}=8.0$
The average solution is given by $\hat{x}=\pi^{1}x^{1}+\pi^{2}x^{2}$, resulting in $\{5,7\}$. Due to the simplicity of the problem, both the (1) and Algorithm 2 managed to yield a solution at the root node that satisfies both the non-anticipativity and integrality conditions ($x_{1}^{1}=x_{1}^{2}=5.0$ and $x_{2}^{1}=x_{2}^{2}=7.0$). However, in the general case, there is no guarantee that this will occur. Therefore, for the sake of illustration, let us assume that the solution to Problem (46) violates one of the conditions, considering $\pi^{1}=\pi^{2}=0.5$.
First, let us assume that $x_{1}^{1}=4$, $x_{1}^{2}=5$, and $x_{2}^{1}=x_{2}^{2}=7$. In this case, the average solution is $\hat{x}=\{4.5,7\}$. We can observe that the first component takes the value 4.5, which is not integer-valued. In this situation, following Algorithm 4, we would perform branching based on the violation of the integrality condition. Specifically, we would generate two child nodes $N_{1}$ and $N_{2}$, such that $N_{1}$ contains Problem (46) with the additional constraints $\{x_{1}^{1},x_{1}^{2}\leq 4\}$, while $N_{2}$ contains Problem (46) with the constraints $\{x_{1}^{1},x_{1}^{2}\geq 5\}$.
Alternatively, let us assume that $x_{1}^{1}=4$, $x_{1}^{2}=6$, and $x_{2}^{1}=x_{2}^{2}=7$. In this case, the average solution is $\hat{x}=\{5,7\}$. We can see that the solution $\hat{x}$ satisfies the integrality condition but does not satisfy the non-anticipativity condition for the first component. Specifically, the dispersion is $\sigma_{1}=x_{1}^{2}-x_{1}^{1}=2$. In this case, following Algorithm 4, we would perform branching based on the violation of the non-anticipativity condition. We would create two child nodes $N_{1}$ and $N_{2}$, where the subproblem for $N_{1}$ corresponds to Problem (46) with the constraints $\{x_{1}^{1},x_{1}^{2}\leq 5-\epsilon_{BB}\}$, and $N_{2}$ would correspond to Problem (46) with the constraints $\{x_{1}^{1},x_{1}^{2}\geq 5+\epsilon_{BB}\}$.